# An explainable three dimensional framework to uncover learning patterns: A unified look in variable sulci recognition

### v3

@@ Abstract
Detecting the significant features of the learning process of an artificial intelligence framework in the entire training and validation dataset can be determined as ’global’ explanations. Studies in the literature lack of accurate, low-complexity, and three-dimensional (3D) global explanations which are crucial in neuroimaging, a field with a complex representational space that demands more than basic two-dimensional interpretations. To fill this gap, we developed a novel explainable artificial intelligence (XAI) 3D-Framework that provides robust, faithful, and low-complexity global explanations. We evaluated our framework on various 3D deep learning networks trained, validated, and tested on a well-annotated cohort of 596 subjects from the TOP-OSLO study. The focus was on the presence and absence of the paracingulate sulcus, a variable feature of brain morphology correlated with psychotic conditions. Our proposed 3D-Framework outperforms traditional XAI methods in terms of faithfulness for global explanations. As a result, we were able to use these robust explanations to uncover new patterns that not only enhance the credibility and reliability of the training process but also reveal promising new biomarkers and significantly related sub-regions. For the first time, our developed 3D-Framework proposes a way for the scientific community to utilize global explanations to discover novel patterns in this specific neuroscientific application and beyond. This study can helps improve the trustworthiness of AI training processes and push the boundaries of our understanding by revealing new patterns in neuroscience and beyond.
@@ 1 Introduction
In both medical imaging and neuroscience, explainability holds paramount importance. A recent study Mamalakis et al. (2024) has categorized the necessity for artificial intelligence (AI) explanations into four categories: self-explainable, semi-explainable, non-explainable applications, and new-pattern discovery. This categorization is based on the variability of expert opinions and the stability of the evaluation protocol.
@@
Specifically in neuroscience AI applications, where there exists significant variability in evaluation protocols and decision-making among experts, the inherent high uncertainty necessitates a greater need for explainability (non-explainable; Mamalakis et al. (2024)). To this end, applications in neuroscience usually need both "local" methods, which provide explanations for each AI prediction separately, and "global" methods, which derive explanations for the decision-making of the AI across the entire dataset. Such variability underscores the importance of thorough investigation and validation of AI model behavior. Even among experienced professionals, knowledge gaps can persist, where AI has the potential to offer insights, and stabilize the validity and key aspects of the protocols Mitchell et al. (2023). This is particularly true for classification tasks where disease evaluation protocols are not yet firmly established (new-patterns discovery; Mamalakis et al. (2024)). For example, the folding of the human brain, established during gestation, is generally consistent among individuals. Post-natal development will affect sulcal metrics but the pattern of sulcation established at birth remains stable throughout life, preserving crucial prenatal neurodevelopmental information Cachia et al. (2016). These conditions makes them very interesting to study, but inter-individual variability create inherent variability in experts’ opinions and uncertainty in the evaluation protocol for tasks related to brain development and folding.
@@
Detecting and annotating sulcal features from brain MRI is hindered by inter-individual variability in sulcation, and manual annotation is both time-consuming and subjective, impeding efficient use of extensive MRI databases. While, to date, automated methods excel in the detection of most sulci, the variable shape and presence/absence of some secondary sulci present a more complex computational hurdle Borne et al. (2020). Successful automation through generalized, unbiased annotation would greatly aid studies focusing on brain folding variations, which are proxies for a critical developmental periods and may relate to cognitive, behavioral, and developmental outcomes, along with psychiatric and neurological disorders. Brain folding is linked to brain function, and specific folding patterns correlate with susceptibility to neurological issues Jiang et al. (2021). For instance, the paracingulate sulcus (PCS), a variable secondary sulcus, is associated with cognitive performance and hallucinations in schizophrenia Fornito et al. (2006),Garrison et al. (2015),Gay et al. (2021),Ćurčić-Blake et al. (2023). To this end, developing networks that can identify the presence or absence of PCS (or multiple PCS elements) in structural MRI and frameworks that can provide both local and global explanations would allow for a more systematic assessment of its functional relevance and impact on critical functions in cognitive, behavioral, and developmental studies Simons et al. (2017).
@@
However, an important challenge arises, particularly in neuroimaging, where the high representation dimensionality in AI applications, the high complexity, the involvement of unstable evaluation protocols, and the intricate expert annotations make these tasks challenging to explain and evaluate. To this end, ordinary 2D local or global XAI methods will fall short and may even be incorrect in some cases Mamalakis et al. (2024). Thus, there is an urgent need for 3D explanation frameworks that can surpass traditional XAI methods and deliver more reliable global explanations, which can be used to validate and identify new patterns in neuroscience.
@@
Contribution. For the first time, we propose a comprehensive framework to validate and explain deep learning models in a way that reveals novel insights into pattern learning, thereby enhancing the credibility and reliability of these networks. To achieve this, we developed a novel XAI 3D-Framework, designed to assess the ability of deep learning networks to learn patterns in a binary classification task to detect the paracingulate sulcus, a highly variable secondary sulcus related to reality monitoring and psychotic conditions.
@@
We applied two distinct local XAI methods, GradCam and SHAP, in 3D space to the entire dataset used in the binary classification task. A dimensionality reduction technique was then employed to extract global explanations from these XAI methods, as well as statistical patterns from the 3D brain inputs. By combining these two XAI methods with the statistical patterns in various configurations, we outperformed traditional XAI methods in global explainability, while also revealing the learning patterns within the network.
@@
For the prediction of the paracingulate sulcus presence, a complex and high uncertainty computational hurdle, we developed two different 3D deep learning networks: a simple 3D convolutional neural network (simple-3D-CNN) and a two-headed attention layer network with variations in backbone options (2CNN-3D-MHL and simple-3D-MHL). The networks utilized 3D brain inputs derived from preprocessing structural MRI scans, which included grey-white surfaces and sulcal skeletons pre-processing images from both the right and left hemispheres. We used a well-annotated cohort of 596 subjects from the TOP-OSLO study Mørch-Johnsen et al. (2017) to train, validate, and test our deep learning networks, with a split of 70% for training, 20% for validation, and 10% for testing.
@@
The proposed 3D framework surpasses traditional XAI methods in terms of faithfulness for global explanation. Importantly, our framework effectively identifies significant sub-regions by combining explainable features with statistical characteristics from the evaluation dataset. This approach offers a robust means to determine the patterns that contribute to a network’s decisions, thereby providing greater transparency, trust and further exploration in the outcomes of deep learning models used in neuroscience applications.
@@ 2.1 The application: Sulcal pattern studies
Cortical folding, which almost entirely develops in utero, results in an inter-individual variability that is often eschewed in populational studies. Yet, focusing on the variability of sulcal pattern can be motivated by different reasons: strict descriptive anatomy, refinement of inter-subject registration, investigation of mechanisms at play during neurodevelopment, and search for anatomo-functional correlates. Links to brain function can also be addressed to either investigate healthy functional variability (e.g. cingulate folding pattern and functional connectivity Fedeli et al. (2022)), or relationships between cortical folding pattern and pathological outcomes (e.g. paracingulate folding and hallucinations in schizophrenia Rollins et al. (2020)). The study of cortical folding variability can be addressed through global methods, with whole brain or regional considerations of generic sulcal parameters (e.g. gyrification index or sulcal pits), but most refined investigations of sulcal patterns require the capability for a focus on specific sulci, hence the need for automated sulcal recognition methods. A number of different methods have been developed for automated sulcal labelling in a general setting (for review, see Mangin et al. (2015)). Yet, to the best of our knowledge, no method is currently implemented to automatically label the paracingulate sulcus in a 3 dimensional approach (a 2 dimensional approach for a specific MRI slice has been proposed Yang et al. (2019)). The fact that the paracingulate sulcus is generally ommited in whole-brain labelling is probably related to the way in which it is anatomically specified: it is not only defined by its location in the brain but also by its orientation (parallel to the cingulate sulcus). Hence, even newer sulcal labelling implementations do not identify it within the folding of the medial frontal cortex Borne et al. (2020). Nevertheless, investigating the whole-brain anatomical correlates of sulcal variability incurred by the presence of the paracingulate sulcus is potentially important to understand some of the common and severe psychotic symptoms. Automatic detection of the paracingulate sulcus is in itself useful for large dataset exploration, especially since it has previously been related to functional variability in reality monitoring Simons et al. (2017). Additionally and unprecedentedly, adding an explainability component allows us to interrogate the anatomical covariates of the paracingulate sulcus gaining a borader vision on the mechanisms at which it may have a role.
The automated labelling of sulci allow shape analysis and pattern investigation of specific sulci on large datasets. These analyses can be led using different methods, on which the following are a general overview. First, simple metrics can be extracted automatically using the BrainVISA software Cointepas et al. (2001), such as the length or mean depth Mangin et al. (2020). Similarily, the local gyrification index can be derived from these automatically extracted metrics, defined as the ratio between a sulcus’s surface area and the outer cortical area Cachia et al. (2008). Yet, these metrics provide limited information on sulcal pattern. Automated methods providing more advanced sulcal pattern characterization have been proposed, as, for example, the comparison of positioning of the superior temporal sulcus between two cohorts Plaze et al. (2011). Another way of investigating sulcal pattern is by automatically capturing shape features. This can be done by looking for a predefined shape feature, such as a characteristic bulk in a sulcus Coulon et al. (2015) or a sulcus’s depth profile Leroy et al. (2015). Alternatively, this can also be investigated by using machine learning to automatically extract shape features from either a sulcus or a group of sulci Sun et al. (2016); De Vareilles et al. (2023); de Vareilles et al. (2023); Meng et al. (2020), with no prior expectation on the shape feature to capture. In addition to offering methodologies for investigating sulcal shape variability, these studies highlight our limited current understanding of the implications of sulcal pattern variability.
@@ 2.2 Explainable methods
Recent studies have seen a significant surge in the exploration of XAI within medical image analysis and neuroimaging domains Samek et al. (2021); van der Velden et al. (2022); Quellec et al. (2021); Mamalakis et al. (2021). XAI methodologies are broadly categorized into interpretable and post-hoc approaches. Interpretable methods focus on models that possess inherent properties such as simulatability, decomposability, and transparency, often linked to linear techniques like Bayesian classifiers, support vector machines, decision trees, and K-nearest neighbor algorithms van der Velden (2023). On the other hand, post-hoc methods are typically used with AI techniques to reveal nonlinear mappings within complex datasets Samek et al. (2021); van der Velden (2023).
A widely used post-hoc technique is Local Interpretable Model-Agnostic Explanations (LIME), which explains the network’s predictions by building simple interpretable models that approximate the deep network locally, i.e., in the close neighborhood of the prediction of interest Singh et al. (2020). Post-hoc techniques include both model-specific approaches that address specific nonlinear behaviors and model-agnostic approaches that explore data complexity Samek et al. (2021); van der Velden (2023). In computer vision, model-agnostic methods such as LIME and perturbation techniques are widely used, while model-specific methods encompass feature relevance, condition-based explanations, and rule-based learning Bach et al. (2015); Rajani et al. (2019); van der Velden (2023).
In medical imaging, explainable methods often focus on attribution and perturbation techniques Tjoa and Guan (2020). Attribution techniques like LIME, LRP, GRAD-CAM, and SHAP, they identify important features for a given prediction by assigning relevance scores to the input features. Perturbation techniques assess the sensitivity of an AI prediction to specific input features by systematically altering sub-groups of the input data Tjoa and Guan (2020); Singh et al. (2020). GRAD-CAM is notably prevalent among explainable methods in medical imaging due to its ease of application and understanding, as well as its ability to map significant features in the imaging space using the activations of the last convolutional layers Singh et al. (2020).
By advancing these explainable methodologies, we can better interpret complex models, enhancing their transparency and trustworthiness, particularly in critical applications like medical imaging and neuroimaging. An important gap in the literature is the absence of three-dimensional representation frameworks that can explain complex models, such as those used in neuroscience, by providing faithful global explanations. Such frameworks could offer more accurate interpretations than established approaches, potentially improving AI transparency and uncovering new patterns in significant sub-regions involved in predictions.
@@ 3 Methods and Background
In this section, we describe the different 3D classifiers used to address the binary classification problem of determining the presence or absence of PCS. Furthermore, we developed a novel XAI 3D-Framework for the non-explainable and new-patterns discovery tasks (Mamalakis et al. (2024)). Specifically, we utilized two distinct explanation methods, belonging to the post-hoc family, SHAP and GradCam, expanded in three-dimensional space. We concatenate the outcomes obtained from the two explainable methods with the results derived from a statistical feature extraction model. The statistical model was a transparent dimensionality reduction algorithm applied to the inputs data of the validation cohort. This whole approach was intended to mitigate any potential biases and enhance the robustness of the pattern learning discovery.
@@ 3.1 Architectural design of deep learning networks for the classification task
We used two different deep learning model for the binary classification task (Fig. 1b.,c.). The first network was a 3D Convolutional Neural Network (CNN) with five levels (see Fig. 1b.). Each level incorporated a CNN block with a 3D convolution layer, a 3D max-pooling layer, and a batch normalization layer. In the first three levels, the 3D convolution layer employed 64, 128, and 256 filters, respectively (as shown in Fig. 1b.). The last level connected to a multi-layer perceptron (MLP) for the final prediction (as illustrated in Fig. 1a.). The MLP comprised three distinct perceptrons and two dropout layers to estimate epistemic uncertainty.
@@
Figure 1: The input modalities and the architecture of simple-3D-CNN and the simple-3D-MHL networks. a, Input modalities illustrated on a right hemisphere coronal slice. More analytically the raw MRI of a given subject, the corresponding grey-white surface, and the corresponding sulcal skeleton. b, The architecture of simple-3D-CNN network with explanation of 3D Convolution layer (3D Conv) and 3D Max Pooling layer on the left. c, The three dimension MHL model with two different backbone choices, the full simple-3D-CNN (simple-3D-MHL) and the two level simple-3D-CNN layer (2CNN-3D-MHL).
For the second network (Fig. 1c.), we used a combination of multi-head attention layers (MHL) to focus on the global diversity and variation of a backbone output. To reduced the biased choice of only one backbone selection, we opted for two distinct backbone networks: (i) a 3D Convolution layer block with two level of 64 and 128 filters a Global Average Pooling, and a perceptron of 32400 hidden layers (2CNN-3D-MHL), and (ii) the straightforward 3D CNN network outlined previously. (Fig. 1b., simple-3D-MHL). We used the multi-head attention mechanism described in Vaswani et al. (2017) and as we used a two-head attention of the same input, we called this approach "self-attention". The two heads ($head_{c}$) are given by:
@@
where $c$ is the backbone output. There are two outputs as we used the backbone twice, and each output is connected to a perceptron with $N$ hidden layers (where $N$ is equal to the product of the weight and height of the input image). The $AT$ is the attention layer and it computed by:
@@
where the input matrix are combination of queries and keys of dimension $d_{k}$, and values of dimension $d_{v}$. Queries are packed together into matrix $Q$. The keys and values are also packed together into matrices $K$ and $V$. The output of the MHL network is given from:

### v4

@@ Abstract
The significant features identified in a representative subset of the dataset during the learning process of an artificial intelligence model are referred to as a ’global’ explanation. Three-dimensional (3D) global explanations are crucial in neuroimaging where a complex representational space demands more than basic two-dimensional interpretations. Curently, studies in the literature lack accurate, low-complexity, and 3D global explanations in neuroimaging and beyond. To fill this gap, we develop a novel explainable artificial intelligence (XAI) 3D-Framework that provides robust, faithful, and low-complexity global explanations. We evaluated our framework on various 3D deep learning networks trained, validated, and tested on a well-annotated cohort of 596 MRI images. The focus of detection was on the presence or absence of the paracingulate sulcus, a highly variable feature of brain topology associated with symptoms of psychosis. Our proposed 3D-Framework outperformed traditional XAI methods in terms of faithfulness for global explanations. As a result, these explanations uncovered new patterns that not only enhance the credibility and reliability of the training process but also reveal the broader developmental landscape of the human cortex. Our XAI 3D-Framework proposes for the first time, a way to utilize global explanations to discover the context in which detection of specific features are embedded, opening our understanding of normative brain development and atypical trajectories that can lead to the emergence of mental illness.
@@ 1 Introduction
In both medical imaging and neuroscience, explainability holds paramount importance. Recently, the study by Mamalakis et al. (2024) introduced the necessity of explanations in artificial intelligence (AI) healthcare applications, categorizing them into four types: self-explainable, semi-explainable, non-explainable applications, and new-pattern discovery. This categorization is based on the variability of expert opinions, the stability of the evaluation protocol, and the dimensionality of the problem.
@@
In neuroscience AI applications, there commonly exists significant variability in evaluation protocols and decision-making among experts. The inherent high uncertainty creates a greater need for explainability (as the framework falls in the non-explainable category; Mamalakis et al. (2024)). To this end, applications usually need both "local" methods, which provide explanations for each AI prediction separately, and "global" methods, which derive explanations for the decision-making of the AI across the entire dataset. Such variability underscores the importance of thorough investigation and validation of AI model behavior. Even among experienced professionals, knowledge gaps can persist, and this is where AI has the potential to offer insights and stabilize the validity of key aspects of the evaluation protocols (Mitchell et al. (2023)). This is particularly true for classification tasks where the key features of a disease are not yet firmly established (matching the new-patterns discovery category; Mamalakis et al. (2024)).
The folding of the human cortical surface occurs during the perinatal period and remains constant for an individual for the rest of their life, much like a fingerprint, preserving early neurodevelopmental information (Cachia et al. (2016)). Broadly, human brains share many features of cortical folding (sulci), however strong inter-individual variability creates inherent divergence in experts’ opinions when it comes to labelling the more variable sulci, impeding the effort to automatically label sulci in the regions showing most variability.
@@
While, to date, automated methods excel in the detection of most sulci, the variable shape and presence/absence of some sulci present a more complex computational hurdle (Borne et al. (2020)). Successful automation through generalized, unbiased annotation would greatly aid studies focusing on brain folding variations, which are proxies for a critical developmental period with information that may relate to cognitive, behavioral, and developmental outcomes, along with psychiatric and neurological disorders. Brain folding is linked to brain function, and specific folding patterns correlate with susceptibility to neurological issues (Jiang et al. (2021)). In particular, the paracingulate sulcus (PCS), a highly variable sulcus, is associated with cognitive performance and hallucinations in schizophrenia (Fornito et al. (2006), Garrison et al. (2015), Gay et al. (2021), Ćurčić-Blake et al. (2023)). To this end, developing networks that can identify the presence or absence of PCS (or multiple PCS elements) in brain magnetic resonance imaging (MRI) and creating frameworks that can provide both local and global explanations would allow for a more systematic characterization and a more comprehensive understanding of whole-brain correlates related to its presence. This, in turn, would allow for both more precise and holistic assessments of its functional relevance and impact on brain functions (Simons et al. (2017)).
@@
However, an important challenge arises, particularly in neuroimaging, where the high representation dimensionality in AI applications, the high anatomical complexity, the involvement of unstable evaluation protocols, and the intricate yet uncertain expert annotations make these tasks challenging to explain and evaluate. Ordinary 2D local or global XAI methods fall short and may even be incorrect in some cases (Mamalakis et al. (2024)). Thus, there is a need for 3D explanation frameworks that can surpass traditional XAI methods and deliver more reliable global explanations.
@@
Contribution. For the first time, we propose a comprehensive framework to validate and explain deep learning models in a way that reveals novel insights into pattern learning, thereby enhancing the credibility and reliability of these networks. To achieve this, we developed a novel XAI 3D-Framework designed to assess the ability of deep learning networks to learn patterns in a binary classification task to detect the presence or absence of the PCS, a highly variable sulcus related to reality monitoring and psychotic conditions.
@@
We applied two distinct local XAI methods, GradCam and SHAP, in 3D space to the entire dataset used in the binary classification task. A dimensionality reduction technique was then employed to extract global explanations from these XAI methods, as well as statistical patterns from the 3D brain inputs. By combining these two XAI methods with the statistical patterns in various configurations we outperformed traditional XAI methods in global explainability while revealing the learning patterns within the network.
@@
For the detection of the PCS presence we developed two different 3D deep learning networks: a simple 3D convolutional neural network (simple-3D-CNN) and a two-headed attention layer network with variations in backbone options (2CNN-3D-MHL and simple-3D-MHL). The networks utilized 3D brain inputs derived from preprocessing structural MRI scans, which included grey-white surface boundaries and sulcal skeletons from both the right and left hemispheres. We used a well-annotated cohort of 596 subjects from the TOP-OSLO study cohort (Mørch-Johnsen et al. (2017)) to train, validate, and test our deep learning networks, with a split of 70% for training, 20% for validation, and 10% for testing.
@@
The proposed 3D framework surpassed traditional XAI methods in terms of faithfulness for global explanation. Importantly, our framework effectively identified significant sub-regions by combining explainable features with statistical characteristics from the evaluation dataset. This approach offers a robust means to determine the patterns that contribute to a network’s decisions, thereby providing greater transparency, trust and further exploration of the outcomes of deep learning models used in neuroscience applications.
@@ 2.1 The application: Sulcal pattern studies
Cortical folding, which develops in the perinatal period (i.e. in the last few months of gestation and the first few months after birth) results in an inter-individual variability that is often eschewed in populational studies. Interest in the variability of sulcal patterns can be motivated by different reasons: strict descriptive anatomy, refinement of inter-subject registration, investigation of mechanisms at play during neurodevelopment, and a search for anatomo-functional correlates. Links to brain function can also be addressed to investigate healthy functional variability (e.g. cingulate folding pattern and functional connectivity Fedeli et al. (2022)), or relationships between cortical folding pattern and pathological outcomes (e.g. paracingulate folding and hallucinations in schizophrenia Rollins et al. (2020)). The study of cortical folding variability can be addressed through global methods, with whole brain or regional considerations of generic sulcal parameters (e.g. gyrification index or sulcal pits), but most refined investigations of sulcal patterns require the capability for a focus on specific sulci, hence the need for automated sulcal recognition methods. A number of different methods have been developed for automated sulcal labelling in a general setting (for review, see Mangin et al. (2015)). Yet, to the best of our knowledge, no method is currently implemented to automatically label the PCS in a 3D approach (a 2D approach for a specific MRI slice has been proposed Yang et al. (2019)). The fact that the PCS is generally omitted in whole-brain labelling is related to the way in which it is anatomically specified; not only by its location in the brain, but also by its orientation (parallel to the cingulate sulcus). Hence, even newer sulcal labelling implementations do not identify the PCS within the folding of the medial frontal cortex (Borne et al. (2020)). Nevertheless, investigating the whole-brain anatomical correlates of sulcal variability incurred by the presence of the PCS is potentially important to understand some of the most common and severe symptoms of psychosis. Automatic detection of the PCS is in itself useful for large dataset exploration, especially since it has previously been related to functional variability in reality monitoring (Simons et al. (2017)). Additionally and unprecedentedly, adding an explainability component allows us to interrogate the anatomical covariates of the PCS gaining a broader vision of the functional mechanisms in which it may have a role.
@@ 2.2 Explainable methods
Recent studies have seen a significant surge in the exploration of XAI within medical image analysis and neuroimaging domains (Samek et al. (2021); van der Velden et al. (2022); Quellec et al. (2021); Mamalakis et al. (2021)). XAI methodologies are broadly categorized into interpretable and post-hoc approaches. Interpretable methods focus on models that possess inherent properties such as simulatability, decomposability, and transparency, often linked to linear techniques like Bayesian classifiers, support vector machines, decision trees, and K-nearest neighbor algorithms (van der Velden (2023)). On the other hand, post-hoc methods are typically used with AI techniques to reveal nonlinear mappings within complex datasets (Samek et al. (2021); van der Velden (2023)).
A widely used post-hoc technique is Local Interpretable Model-Agnostic Explanations (LIME), which explains the network’s predictions by building simple interpretable models that approximate the deep network locally, i.e. in the close neighborhood of the detected structure (Singh et al. (2020)). Post-hoc techniques include both model-specific approaches that address specific nonlinear behaviors and model-agnostic approaches that explore data complexity (Samek et al. (2021); van der Velden (2023)). In computer vision, model-agnostic methods such as LIME and perturbation techniques are widely used, while model-specific methods encompass feature relevance, condition-based explanations, and rule-based learning (Bach et al. (2015); Rajani et al. (2019); van der Velden (2023)).
In medical imaging, explainable methods often focus on attribution and perturbation techniques (Tjoa and Guan (2020)). Attribution techniques like LIME as well as Layer-wise Relevance Propagation (LRP), Gradient-weighted Class Activation Mapping (GradCAM), and Shapley Additive Explanations (SHAP) identify important features for a given prediction by assigning relevance scores to the input features. Perturbation techniques assess the sensitivity of an AI prediction to specific input features by systematically altering sub-groups of the input data (Tjoa and Guan (2020); Singh et al. (2020)). GradCAM is notably prevalent among explainable methods in medical imaging due to its ease of application and understanding, as well as its ability to map significant features in the imaging space using the activations of the last convolutional layers (Singh et al. (2020)).
By advancing these explainable methodologies we can better interpret complex models, enhancing their transparency and trustworthiness, particularly in applications like medical imaging, and neuroimaging. An important gap in the literature is the absence of three-dimensional representation frameworks that can explain complex models, such as those used in neuroscience, by providing faithful global explanations. Such frameworks could offer more accurate interpretations than established approaches, potentially improving AI transparency and uncovering new patterns in significant sub-regions involved in classification and prediction.
@@ 3 Methods and Background
In this section, we describe the different 3D classifiers used to address the binary classification problem of determining the presence or absence of PCS. Furthermore, we develop a novel XAI 3D-Framework for the non-explainable and new-patterns discovery tasks (Mamalakis et al. (2024)). We utilize two distinct explanation methods, belonging to the post-hoc family SHAP and GradCam, expanded in three-dimensional space. We concatenate the outcomes obtained from the two explainable methods with the results derived from a statistical feature extraction model. The statistical model is a transparent dimensionality reduction algorithm applied to the input data of the validation cohort. This whole approach is intended to mitigate any potential biases and enhance the robustness of the pattern learning discovery.
@@ 3.1 Architectural design of deep learning networks for the classification task
We used two different deep learning models for the binary classification task (Fig. 1b.,c.). The first network was a 3D Convolutional Neural Network (CNN) with five levels (see Fig. 1b.). Each level incorporated a CNN block with a 3D convolution layer, a 3D max-pooling layer, and a batch normalization layer. In the first three levels, the 3D convolution layer employed 64, 128, and 256 filters, respectively (as shown in Fig. 1b.). The last level connected to a multi-layer perceptron (MLP) for the final prediction (as illustrated in Fig. 1a.). The MLP comprised three distinct perceptrons and two dropout layers to estimate epistemic uncertainty.
@@
Figure 1: The input modalities and the architecture of simple-3D-CNN and the simple-3D-MHL networks. a, Input modalities illustrated on a right hemisphere coronal slice. In more detail, the raw MRI of a given subject, the corresponding grey-white surface, and the corresponding sulcal skeleton. b, The architecture of simple-3D-CNN network with explanation of 3D Convolution layer (3D Conv) and 3D Max Pooling layer on the left. c, The three dimension MHL model with two different backbone choices, the full simple-3D-CNN (simple-3D-MHL) and the two level simple-3D-CNN layer (2CNN-3D-MHL).
For the second network (Fig. 1c.), we used a combination of multi-head attention layers (MHL) to focus on the global diversity and variation of a backbone output. To reduce the biased choice of only one backbone selection, we opted for two distinct backbone networks: (i) a 3D Convolution layer block with two levels of 64 and 128 filters a Global Average Pooling, and a perceptron of 32400 hidden layers (2CNN-3D-MHL), and (ii) the straightforward 3D CNN network outlined previously (Fig. 1b., simple-3D-MHL). We used the multi-head attention mechanism described in (Vaswani et al. (2017)) and as we used a two-head attention of the same input ("self-attention"). The two heads ($head_{c}$) are given by:
@@
where $c$ is the backbone output. There are two outputs corresponding to the two uses of the backbone, and each output is connected to a perceptron with $N$ hidden layers (where $N$ is equal to the product of the weight and height of the input image). The $AT$ is the attention layer and it computed by:
@@
where the input matrix are combinations of queries and keys of dimension $d_{k}$, and values of dimension $d_{v}$. Queries are packed together into a matrix $Q$. The keys and values are also packed together into matrices $K$ and $V$. The output of the MHL network is given by: